OpenStudy (anonymous):

How do you write x^4+12x^2-8 in quadraic form

OpenStudy (amistre64):

take a square out of the exponents

OpenStudy (amistre64):

(x^2)^2 + 12(x)^2 - 8

OpenStudy (amistre64):

almost wrote it right lol

OpenStudy (amistre64):

say u = x^2; then u^2 = x^4 u^2 +12u -8

OpenStudy (anonymous):

Thank-you!